# Exploring the Small World of Word Embeddings: A Comparative Study on Conceptual Spaces from LLMs of Different Scales

### v1

@@ 4.4 Scenario 3: SMM of Qualitative Words
In Scenario 3, we utilize a cross-lingual semantic map for adjectives and qualitative words Perrin (2010), which includes 22 African languages, French, and English. Polysemes in each language are connected to indicate conceptual proximity. The domain spans dimension, age, value, color, etc., with capitalized English words representing concepts, such as BIG, SMALL, LONG, SHORT, WIDE, and DEEP for dimension. We filter out words not present in the Llama vocabulary, resulting in 75 concepts. The final set of concepts and the human-annotated graph are presented in Appendix A.2.
@@ 5 Results and Analysis
In this section, we first construct the conceptual space based on the minimum connectivity approach described in Section 3.3. We then evaluate the space in three distinct scenarios.
@@ 5.1 Graph Construction
ASPL ($\downarrow$) 3.392 3.353
Table 2: Statistics for Llama2-7B and Llama2-70B. In the “Small-world” section, GCC refers to global clustering coefficient, ALCC to average local clustering coefficient, and Diameter and ASPL refer to the longest and average shortest path lengths, respectively.
@@
We present the statistics of the conceptual spaces for both models in Table 2, divided into three parts. The Basic section includes the number of nodes (#Nodes), edges (#Edges), and the average (Avg. Degree) and standard deviation (Std. Degree) of degrees. The Weighted section calculates the average and standard deviation of weighted degrees (also called Strength or Traffic), along with the equivalent threshold—the minimum weight value in the final graph $G$. The last section focuses on Small-world effects. The global clustering coefficient (GCC) measures graph transitivity—the fraction of actual triangles among all possible triangles in $G$. The local clustering coefficient (ALCC) is calculated as the average of actual connections within neighbors for all nodes. The network diameter is the longest path between any two nodes. Smaller diameters and larger GCC/ALCC values indicate stronger small-world effects. We also calculate the average shortest path length (ASPL) for both models.
@@
Both models exhibit strong small-world clustering, with high GCC and ALCC values and a short diameter. In contrast, random networks with the same edge count have much lower GCC (0.0032) and ALCC (0.0020). The observed diameter of 6 aligns with the Six Degrees of Separation theory Milgram (1967), commonly found in social networks Watts and Strogatz (1998) and web structures Albert et al. (1999).
@@ 5.2 Scenario 1
In this scenario, we evaluate the conceptual spaces of both models by calculating the shortest path length between any pair of nodes within and between semantic groups. The shortest path is computed using Dijkstra’s algorithm from the NetworkX package  [^3] https://networkx.org/documentation/stable/reference/algorithms/shortest_paths.html . The average lengths for Llama2-7B and Llama2-70B are shown in Figure 3. Figure 4 presents the difference heatmap, where the difference is computed as the average length of Llama2-70B minus that of Llama2-7B. The significance of the differences is indicated by the number of stars (one, two, or three), corresponding to p-values of 0.05, 0.01, and 0.001 in the t-test, respectively.
Our results show that Llama2-70B generally has longer path lengths than Llama2-7B, both within and between semantic groups, except in the RANDOM vs. NAME or PLACE comparison. This suggests that Llama2-70B follows more complex paths to uncover relationships. Figures 5 and 6 illustrate this with the six shortest paths from “bar” to “library.” The shortest path (red) in Llama2-70B shows greater variation in word form and conjugation. Additionally, its other paths include multilingual instances and diverse connections, aiding disambiguation—for example, distinguishing “library” as a building from its software-related meaning.
Figure 5: Paths between “bar” and “library” for Llama2-7B. Edge width reflects weight values. Nodes and edges along the shortest path in terms of the summed weights are marked in red.
Figure 6: Paths between “bar” and “library” for Llama2-70B. Edge width represents weight values. Nodes and edges in the shortest path are highlighted in red. The paths illustrate a complex relationship involving similar word forms, multilingual links, and disambiguation groups.
Figure 7: Paths between “collection” and “set” in Llama2-7B. The red line represents the shortest path in terms of the summed weights.
Figure 8: Paths between “collection” and “set” in Llama2-70B. The red line indicates the shortest path, which follows a more logical transition from a singular concept to a plural one.
@@ 5.3 Scenario 2
The results are similar to Scenario 1: Llama2-70B generally has longer paths than Llama2-7B. For most relations, the length of 70B is significantly longer, except for those related to word form, where both models have direct edges for the word pair. Longer paths appear in member of domain topic, verb group, and hypernym relations, indicating a more complex connection chain. For example, in the hypernym relation between “set” and “collection”, Llama2-70B shows a more intricate connection, gradually linking the singular “set” to the plural “collection” with some conjugation correlation, as seen in Figures 7 and 8.
@@ 5.4 Scenario 3
In Scenario 3, we construct a conceptual space for qualitative words, with a reference (GT) from cross-lingual research. The statistics for the spaces built by Llama2-7B (7B), Llama2-70B (70B), and GT are shown in Table 3.
@@
In general, the spaces constructed by the models have more edges than those built by human experts, with fewer connected components and isolated nodes. This is because experts create the graph by collecting a corpus and adding an edge only when at least three languages exhibit concept co-occurrence. This process is especially challenging for low-resource languages. In contrast, the vectorized concepts from models generate a much denser graph. Furthermore, Llama2-70B tends to be sparser than Llama2-7B. Compared to GT, the automatic space aligns well, showing moderate correlation and coverage (indicated by recall). However, the precision is lower due to the larger number of edges, suggesting that embeddings could be used to initially construct a space, which linguists could later refine.
@@ 6 Conclusion
This paper investigates the construction of conceptual spaces using input embeddings from large language models (LLMs). We analyze and compare the network properties of two LLMs with different scales across three scenarios. Our findings show that conceptual spaces can be effectively constructed from embeddings, which exhibit a small-world clustering effect. Additionally, models with more parameters tend to explore longer and more complex paths between concepts, partially supporting the “scaling law” Kaplan et al. (2020). This study also provides an efficient approach to constructing conceptual spaces, potentially benefiting fields such as language typology and cognitive science.
@@ 7 Limitations
We acknowledge several limitations in our work. First, we represent each concept as a word without considering context. However, words can be ambiguous, particularly for homonyms that encompass multiple unrelated meanings. Second, our evaluation is limited to two models, Llama2-7B and Llama2-70B. Results may differ with models of different architectures or parameter scales. Additionally, we focus only on input embeddings and do not explore the properties of output embeddings, which may also capture individual word representations. Finally, the length of the shortest path in conceptual spaces is not a definitive metric for embedding quality, and we plan to explore more sophisticated metrics to better reflect the characteristics of these spaces.
@@ Appendix A Appendix
A.1 Common Concepts in Scenario 1
@@
Table 4: Specific concepts from different semantic groups.
A.2 Conceptual Spaces in Scenario 3

### v2

@@ 4.4 Scenario 3: SMM of Qualitative Words
In Scenario 3, we leverage a cross-linguistic semantic map of adjectives and qualitative terms Perrin (2010), which encompasses 22 African languages along with French and English. In this resource, polysemous words within each language are linked to reflect conceptual proximity, forming a semantic map that encodes a universal network of meanings. This map captures cross-linguistic regularities in polysemy patterns, serving as an invariant conceptual structure underlying diverse lexical realizations. The semantic domains covered include dimension, age, value, and color, with concepts represented by capitalized English words (e.g., BIG, SMALL, LONG, SHORT, WIDE, DEEP for the dimension domain).
We filter out words not present in the LLaMA vocabulary, resulting in a final set of 75 concepts from the original 110. The full list of concepts and the human-annotated semantic graph are provided in Appendix A.3. This scenario evaluates how well the embedding-induced lexico-semantic network aligns with a cross-linguistically grounded conceptual structure.
@@ 5 Results and Analysis
In this section, we first construct the lexico-semantic network based on the minimum connectivity approach described in Section 3.3. We then evaluate the space in three distinct scenarios.
@@ 5.1 Graph Construction
Table 2: Statistics for Llama2-7B and Llama2-70B. GCC refers to global clustering coefficient, ALCC to average local clustering coefficient, and Diameter refers to the longest path length. A star () indicates high statistical significance between the two models for that metric. Metrics without a star are not tested for significance, typically because they are fixed by design (e.g., #Nodes, #Edges) or not applicable for statistical comparison.
@@
We present the statistics of the lexico-semantic networks for both models in Table 2, divided into three parts. The Basic section includes the number of nodes (#Nodes), edges (#Edges), and the average (Avg. Degree) and standard deviation (Std. Degree) of degrees. The Weighted section calculates the average and standard deviation of weighted degrees (also called Strength or Traffic), along with the equivalent threshold—the minimum weight value in the final graph $G$. The last section focuses on Small-world effects. The global clustering coefficient (GCC) measures graph transitivity—the fraction of actual triangles among all possible triangles in $G$. The local clustering coefficient (ALCC) is calculated as the average of actual connections within neighbors for all nodes. The network diameter is the longest path between any two nodes. Smaller diameters and larger GCC/ALCC values indicate stronger small-world effects. We also calculate the statistical difference of t-test and mark it by a star (*) for a high difference (with a p value less than 0.05).
@@
Both models exhibit pronounced small-world characteristics, reflected in their high global clustering coefficients (GCC), average local clustering coefficients (ALCC), and short diameters. In contrast, random networks with the same number of edges show substantially lower GCC (0.0032) and ALCC (0.0020). Notably, the observed diameter of 6 corresponds with the Six Degrees of Separation theory Milgram (1967), a phenomenon commonly observed in social networks Watts and Strogatz (1998) and web graphs Albert et al. (1999). Compared to the smaller model, the 70B model displays a somewhat weaker small-world effect, evidenced by higher GCC and ALCC values, suggesting longer and more complex node interactions. Our subsequent local analyses provide further support for these observations through concrete examples.
@@ 5.2 Scenario 1
In this scenario, we evaluate the lexico-semantic networks of both models by calculating the shortest path length between any pair of nodes within and between semantic groups. The shortest path is computed using Dijkstra’s algorithm from the NetworkX package  [^3] https://networkx.org/documentation/stable/reference/algorithms/shortest_paths.html with the cosine distance between any two nodes. The average lengths for Llama2-7B and Llama2-70B are shown in Figure 3. Figure 4 presents the difference heatmap, where the difference is computed as the average length of Llama2-70B minus that of Llama2-7B. The significance of the differences is indicated by the number of stars (one, two, or three), corresponding to p-values of 0.05, 0.01, and 0.001 in the t-test, respectively.
Our results indicate that Llama2-70B generally exhibits longer path lengths than Llama2-7B, both within and across semantic groups, suggesting that Llama2-70B uncovers relationships through more complex pathways. Figures 6(a) and 6(b) illustrate the six shortest paths between “bar” and “library.” The shortest path (highlighted in red) in Llama2-70B shows greater variation in word forms and conjugations. Moreover, alternative paths involve multilingual links and diverse connections that assist in disambiguation—for instance, differentiating the building “library” from its software-related sense. This richer semantic representation may account for the superior performance of larger-scale models.
(a) Llama2-7B
(b) Llama2-70B
Figure 6: Shortest paths between “bar” and “library” in Llama2-7B and Llama2-70B. Edge width indicates weight; red highlights mark the shortest path. The 70B model shows a more complex structure involving similar forms, multilingual links, and disambiguation clusters.
(a) Llama2-7B
(b) Llama2-70B
Figure 7: Shortest paths between “collection” and “set” in Llama2 models. Red highlights indicate the shortest path. The 70B model shows a more logical semantic transition from singular to plural concepts.
@@ 5.3 Scenario 2
The results mirror those of Scenario 1: Llama2-70B generally exhibits longer paths than Llama2-7B. For most relation types, the path lengths in 70B are significantly greater, except for word form–related relations where both models maintain direct edges between word pairs. Longer paths are observed in relations such as member of domain topic, verb group, and hypernym, reflecting more complex connection chains. For instance, in the hypernym relation between “set” and “collection,” Llama2-70B demonstrates a more nuanced path, gradually linking the singular “set” to the plural “collection” with conjugation correlations, as illustrated in Figures 7(a) and 7(b). This suggests that larger-scale models capture more logical and intricate conceptual relations.
@@ 5.4 Scenario 3
In Scenario 3, we construct a lexico-semantic network for qualitative words, with a reference (GT) from cross-lingual research. The statistics for the spaces built by Llama2-7B (7B), Llama2-70B (70B), and GT are shown in Table 3.
@@
Overall, the model-constructed networks contain more edges and fewer connected components and isolated nodes than those built by human experts. Expert-constructed graphs are based on corpus co-occurrence across at least three languages, a process especially difficult for low-resource languages. In contrast, model embeddings produce much denser graphs. Notably, Llama2-70B yields a sparser network than Llama2-7B. Compared to the ground truth (GT), the automatic networks show moderate correlation and coverage (recall), indicating that the lexico-semantic network partially captures universal cross-linguistic patterns of semantic relatedness in polysemy. However, precision is lower due to the larger number of edges, suggesting embeddings can serve as a preliminary space for linguists to further refine.
@@ 6 Conclusion
This paper investigates the construction of lexico-semantic networks using input embeddings from large language models (LLMs). We analyze and compare the network properties of LLMs with different scales across three scenarios. Our findings show that the lexico-semantic network can be effectively constructed from embeddings, which exhibit a small-world clustering effect. Additionally, models with more parameters tend to explore longer and more complex paths between concepts, partially supporting the “scaling law” Kaplan et al. (2020). This study also provides an efficient approach to constructing conceptual spaces, potentially benefiting fields such as language typology and cognitive science.
@@ 7 Limitations
We acknowledge several limitations in our work. First, we mainly concentrate on the monosemous words. However, words can be ambiguous, particularly for homonyms that encompass multiple unrelated meanings. Second, our evaluation is limited to two models, Llama2-7B and Llama2-70B. Results may differ with models of different architectures or parameter scales. Additionally, we focus only on input embeddings and do not explore the properties of output embeddings, which may also capture individual word representations. Finally, the length of the shortest path in the lexico-semantic network is not a definitive metric for embedding quality, and we plan to explore more sophisticated metrics to better reflect the characteristics of these spaces.
@@ Appendix A Appendix
A.1 Construction of Lexico-Semantic Network
Algorithm 1 illustrates the procedure for constructing the lexico-semantic network. The process involves extracting token embeddings from a language model, computing pairwise cosine distances, and retaining a subset of edges to ensure minimal connectivity. This algorithm corresponds to the method described in Section 3.3.
Algorithm 1 Constructing Lexico-Semantic Network with Minimal Connectivity
Token set $\mathcal{V}$, LLM model $M$
Lexico-semantic graph $\mathcal{G}=(V,E)$ and number of selected edges $K$
$V\leftarrow\mathcal{V}$
Extract embeddings $\mathcal{E}$ from $M$ for each $v\in V$
Center each embedding: $\mathcal{E}\leftarrow\mathcal{E}-\text{mean}(\mathcal{E})$
Compute pairwise cosine distances $\text{D}(u,v)$ for all $u,v\in V$
Initialize empty graph $\mathcal{G}=(V,\emptyset)$
Sort all edges $(u,v)$ by increasing $\text{D}(u,v)$ to get edge list $L$
for each edge $(u,v)$ in $L$ do
     Add edge $(u,v)$ to $E$
     if $\mathcal{G}$ is connected then
         return $\mathcal{G},\ K=|E|$
     end if
end for
A.2 Common Concepts in Scenario 1
@@
RANDOM conscious, distance, measure, almost, paste, sun, friend, other, waste, tongue
Table 4: Specific words from different semantic groups.
A.3 Conceptual Spaces in Scenario 3